OpenStudy (anonymous):

If \frac(x^2)(64) + \frac (y^2)(25) = 1 and y( 2 ) = 4.84 , find y'( 2 ) by implicit differentiation.

OpenStudy (anonymous):

Okay, so what do you get when you differentiate the first equation?

OpenStudy (anonymous):

2x/64

OpenStudy (anonymous):

It should still be an equation.

OpenStudy (anonymous):

2x/64+(2yy')/25

OpenStudy (anonymous):

=1

OpenStudy (anonymous):

\[\frac d {dx} 1 = 1 \]Oh?

OpenStudy (anonymous):

no it wasn't

OpenStudy (anonymous):

the problem was changed on me so now its If \frac(x^2)(25) + \frac (y^2)(16) = 1 and y( 1 ) = 3.92 , find y'( 1 ) by implicit differentiation

OpenStudy (anonymous):

Okay, so what is your equation after differentiating.

OpenStudy (anonymous):

2x/25+2yy'/16=0

OpenStudy (anonymous):

now, what is \(x\)?

OpenStudy (anonymous):

You need to plug in the values of \(x\) and \(y\)

OpenStudy (anonymous):

2/25+2(4.84)y'/16=0

OpenStudy (anonymous):

Now solve for \(y'\).

OpenStudy (anonymous):

I'm not going to get a medal, am I?

OpenStudy (mathstudent55):

\( \dfrac{x^2}{25} + \dfrac {y^2}{16} = 1 \) \( \dfrac{2x}{25} + \dfrac{2yy'}{16} = 0 \) \( \dfrac{2(1)}{25} + \dfrac{2(3.92)y'}{16} = 0 \) \( \dfrac{2(3.92)y'}{16} = -\dfrac{2}{25} \) \( y' = -\dfrac{2(16)}{25(2)(3.92) } \) \( y' = -\dfrac{8}{49 } \)